co-efficients of correlation, and The two regression equations from the following informationN= 10, ∑X= 350, ∑Y= 310, ∑(X-35)2 = 162, ∑(Y-31)2 = 222, ∑(X-35)(Y-31)= 92


1
Expert's answer
2021-06-28T04:06:40-0400

"\u2211(x\u221235)^2=162=\u2211dx^2=162"

"\u2211(y\u221231)^2=222=\u2211dy^2=222"

"\u2211(x\u221235(y\u221231)=92=\u2211dxdy=92"

 Regression equation "X- X\u00af = b_{xy} Y-Y\u00af"

"\u200b"

"b_{xy} =\frac {N∑XY - ∑X∑Y}{N∑Y^2 -(∑Y)^2}"


"=\frac{ 10 (108500) -108500}{10(96100) -96100}"


"=\frac{ 976500}{864900}"


"b_{xy}= 1.13"


"x\u00af= \u221a162"

"= 12.73"

"Y\u00af=\u221a222"

"= 14.9"


Regression equation of X on Y


"X - 12.73 = 1.13(Y -14.9)"

"X =1.13Y -2.17"


Regression equation of Y on X

"Y - Y\u00af= b_{xy} (X-X\u00af)"

"= Y - 14.9 = 1.13(X -12.73)""Y = 1.13X + 2.17"